INVESTIGATION OF THE METHYLENE AMIDOGEN RADICAL USING THE CRDS TECHNIQUE

\underline{BORIS NIZAMOV} and PAUL J. DAGDIGIAN, Department of Chemistry, The Johns Hopkins University, Baltimore, Md 21218-2685.

Cavity ringdown spectroscopy (CRDS) was used to study room temperature chemical reactions of the methylene amidogen radical (CH_{2}N). The radical was prepared by 193 nm photolysis of formaldoxime, CH_{2}NOH → CH_{2}N+OH. CRDS signal from both CH_{2}N and OH [A–X (1,0) band] was observed in the wavelength region 278-288 nm. By comparison of the OH and CH_{2}N signals, absorption cross section of the CH_{2}N was determined. To correct for the loss of OH radical by reaction, the rate constant of the reaction of OH with formaldoxime was measured, k(CH_{2}NOH+OH)= 2.5*10^{−12} molecule cm^{3} s^{−1}. Reaction of CH_{2}N with stable molecules such oxygen, hydrogen and methane could not be observed and only an upper limit of the reaction rate constants, < 1.0 *10^{−15} molecule cm^{3} s^{−1}, could be derived. Self-recombination was the main removal process for the CH_{2}N radical under the conditions of our experiment.